-2/11+5x/= -14 / represent absolute value symbol
Answer by stanbon(75887)

-2|11+5x|= -14
----
Divide by -2 to get:
|11+5x| = 7
---
11+5x = 7 OR 11+5x = -7
5x = -4 OR 5x = -18
x = -4/5 OR x = -18/5
================================
Cheers,
Stan H.